OpenStudy (anonymous):

Combine like terms. 8 + 6x – 11x + 12 A. 20 – 5x B. 14x + 1x C. 20 + 5x D. 20 – 17x

RhondaSommer (rhondasommer):

-11 + 6 is what

RhondaSommer (rhondasommer):

welcome to OS! Im Rhonda and would love to help

alones (alones):

Hey Welcome to Openstudy :) Let's simplify step-by-step. 8+6x−11x+12 =8+6x+−11x+12 Combine Like Terms: =8+6x+−11x+12 =(6x+−11x)+(8+12) =−5x+20 I hope it helps :)

RhondaSommer (rhondasommer):

@AloneS i was doin this :(

OpenStudy (anonymous):

thanks

alones (alones):

@RhondaSommer it's fien i'll medal you<3 =]

alones (alones):

no medal for me +'[

OpenStudy (anonymous):

this has to be a cause the answer is 15 and a had 20-5 which equals 15

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct

RhondaSommer (rhondasommer):

can you send me the coding for that @563blackghost

563blackghost (563blackghost):

the greeting?

RhondaSommer (rhondasommer):

yes

563blackghost (563blackghost):

ok ^^